# Quasi-random Multi-Sample Inference for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are often equipped with multi-sample decoding strategies. Vilnis et al. (2023) show that an LLM implicitly defines an arithmetic code book, facilitating efficient and embarrassingly parallelizable arithmetic sampling to produce multiple samples using quasi-random codes. Traditional text generation methods, such as beam search and sampling-based techniques, have notable limitations: they lack parallelizability or diversity of sampled sequences. This study explores the potential of arithmetic sampling, contrasting it with ancestral sampling across two decoding tasks that employ multi-sample inference: chain-of-thought reasoning with self-consistency and machine translation with minimum Bayes risk decoding. Our results demonstrate that arithmetic sampling produces more diverse samples, significantly improving reasoning and translation performance as the sample size increases. We observe a $\mathbf{3\text{-}5\%}$ point increase in accuracy on the GSM8K dataset and a $\mathbf{0.45\text{-}0.89\%}$ point increment in COMET score for WMT19 tasks using arithmetic sampling without any significant computational overhead.
Quasi-random Multi-Sample Inference for Large Language Models
Aditya Parashar thanks: * indicates equal contribution, Aditya Vikram Singh 11footnotemark: 1, Avinash Amballa 11footnotemark: 1, Jinlin Lai, Benjamin Rozonoyer University of Massachusetts, Amherst {aparashar, avsingh,aamballa,jinlinlai,brozonoyer}@umass.edu
@@ 1 Introduction
There have been enormous efforts in improving the performance and efficiency of inference with large language models (Ippolito et al., 2019; Su and Collier, 2023; Grubisic et al., 2024; Zhou et al., 2024; Ding et al., 2024) based on system, data, and model level enhancements. In this paper, we consider that any decoding routine can be broadly assessed by its sample diversity (A.1) and parallelizability. Search-based techniques like beam search can approximate maximum a posteriori (MAP) decoding, mitigating duplicate samples at the expense of not being embarrassingly parallel. Sampling-based methods grounded in ancestral sampling techniques are parallel but don’t explicitly guarantee diverse sequences. The recently proposed arithmetic sampling Vilnis et al. (2023) technique enables parallel inference with diverse samples – by interpreting the inference as sampling from code points from a unit interval, given code points generating sequences becomes embarrassingly parallel and the sample diversity is guaranteed by construction.
Decoding from pre-trained LLMs requires varying strategies for different downstream tasks. For complex reasoning and question-answering tasks, chain-of-thought (CoT) prompting Wei et al. (2022) is established for improving inference by instructing the model to generate intermediate reasoning paths. Wang et al. (2023) propose self-consistency as an additional improvement over chain-of-thought reasoning with multi-sample inference, attributable to diverse reasoning paths enhancing the confidence of the majority answer. For machine translation, minimum Bayes risk (MBR) decoding Kumar and Byrne (2004) is a classical approach for selecting the optimal translation from candidate translations generated by an LLM, requiring diversity to ensure performance.
@@ 2.3 MBR decoding
Minimum Bayes risk (MBR) Eikema and Aziz (2020) is based on the principle of maximizing the expected utility of a given hypothesis. When making predictions, we lack information about the ideal (target) translations and must make decisions under uncertainty. MBR allows the model to probabilistically estimate ideal decisions as it searches for the candidate that maximizes expected utility. We used COMET Rei et al. (2020) as the utility metric.
We use the sampling-based approximation to MBR decoding as posited in Eikema and Aziz (2022), using the Monte Carlo estimate and formulating the candidate space from the generated $N$ (pseudo-reference) samples:
@@ 4 Conclusions
This work investigated the application of arithmetic sampling, a diverse parallelizable sampling technique for multi-sample inference from large language models (LLMs), in chain-of-thought reasoning with self-consistency and machine translation with minimum Bayes risk (MBR) decoding. Experiments on arithmetic reasoning and commonsense reasoning demonstrated that arithmetic sampling improves performance as the number of sampled sequences increases, yielding $3$-$5\%$ and $1.05\%$ point increases in accuracy, respectively, over ancestral sampling at $40$ sampled sequences. In machine translation, arithmetic sampling consistently outperformed ancestral sampling, with improvements ranging from $0.45\%$ to $0.89\%$ in mean COMET scores for $5$ sampled sequences across different language pairs and parameter settings. Arithmetic sampling also exhibited a lower standard deviation compared to ancestral sampling. The superior performance of arithmetic sampling without any significant computational overhead is due to its ability to generate more diverse samples, enhancing LLM performance in tasks like chain-of-thought reasoning and machine translation through parallel and diverse sample generation.

### v2

@@ Abstract
Large language models (LLMs) are often used with decoding strategies that require sampling multiple outputs. Vilnis et al. (2023) show that an LLM implicitly defines an arithmetic code book, facilitating efficient and embarrassingly parallelizable arithmetic sampling to produce multiple samples using quasi-random codes. Traditional text generation methods, such as beam search and sampling-based techniques, have notable limitations: they lack parallelizability or diversity of sampled sequences. This study explores the potential of arithmetic sampling, contrasting it with ancestral sampling across two decoding tasks that employ multi-sample inference: chain-of-thought reasoning with self-consistency and machine translation with minimum Bayes risk decoding. Our results demonstrate that arithmetic sampling produces more diverse samples, significantly improving reasoning and translation performance as the sample size increases. We observe a $\mathbf{3\text{-}5\%}$ point increase in accuracy on the GSM8K dataset and a $\mathbf{0.45\text{-}0.89\%}$ point increment in COMET score for WMT19 tasks using arithmetic sampling without any significant computational overhead.
@@ 1 Introduction
There have been enormous efforts in improving the performance and efficiency of inference with large language models (Ippolito et al., 2019; Su & Collier, 2023; Grubisic et al., 2024; Zhou et al., 2024; Ding et al., 2024) based on system, data, and model level enhancements. In this paper, we consider that any decoding routine can be broadly assessed by its sample diversity (A.1) and parallelizability. Search-based techniques like beam search can approximate maximum a posteriori (MAP) decoding, mitigating duplicate samples at the expense of not being embarrassingly parallel. Sampling-based methods grounded in ancestral sampling techniques are parallel but don’t explicitly guarantee diverse sequences. The recently proposed arithmetic sampling Vilnis et al. (2023) technique enables parallel inference with diverse samples – by interpreting the inference as sampling from code points from a unit interval, given code points generating sequences becomes embarrassingly parallel and the sample diversity is guaranteed by construction.
Decoding from pre-trained LLMs requires varying strategies for different downstream tasks. For complex reasoning and question-answering tasks, chain-of-thought (CoT) prompting Wei et al. (2022) is established for improving inference by instructing the model to generate intermediate reasoning paths. Wang et al. (2023) propose self-consistency as an additional improvement over chain-of-thought reasoning with multi-sample inference, attributable to diverse reasoning paths enhancing the confidence of the majority answer. For machine translation, minimum Bayes risk (MBR) decoding Kumar & Byrne (2004) is a classical approach for selecting the optimal translation from candidate translations generated by an LLM, requiring diversity to ensure performance.
@@ 2.3 MBR decoding
Minimum Bayes risk (MBR) Eikema & Aziz (2020) is based on the principle of maximizing the expected utility of a given hypothesis. When making predictions, we lack information about the ideal (target) translations and must make decisions under uncertainty. MBR allows the model to probabilistically estimate ideal decisions as it searches for the candidate that maximizes expected utility. We used COMET Rei et al. (2020) as the utility metric.
We use the sampling-based approximation to MBR decoding as posited in Eikema & Aziz (2022), using the Monte Carlo estimate and formulating the candidate space from the generated $N$ (pseudo-reference) samples:
@@ 3.3.2 Machine Translation with MBR
3.4 Discussion
We observe that GSM8K demonstrates greater performance gains compared to Commonsense QA and Machine Translation tasks. This is attributed to nature of the tasks. GSM8K involves open-ended reasoning, where there are no predefined options, and diverse reasoning paths can significantly improve performance by exploring a broader solution space. Arithmetic sampling encourages this diversity, enabling the model to better capture complex reasoning patterns.
In contrast, tasks like Commonsense QA or Machine Translation are more constrained. Commonsense QA has only a limited set of options (e.g., 5 choices), reducing the likelihood of errors and the potential impact of diversity. Similarly, in Machine Translation, translations are typically close to the target sentence, leaving little room for diversity to improve outcomes significantly. This explains why arithmetic sampling shows limited gains in these tasks compared to GSM8K.
@@ 4 Conclusions
This work investigated the application of arithmetic sampling, a diverse parallelizable sampling technique for multi-sample inference from large language models (LLMs), in chain-of-thought reasoning with self-consistency and machine translation with minimum Bayes risk (MBR) decoding. Experiments on arithmetic reasoning and commonsense reasoning demonstrated that arithmetic sampling improves performance as the number of sampled sequences increases, yielding $3$-$5\%$ and $1.05\%$ point increases in accuracy, respectively, over ancestral sampling at $40$ sampled sequences. In machine translation, arithmetic sampling consistently outperformed ancestral sampling, with improvements ranging from $0.45\%$ to $0.89\%$ in mean COMET scores for $5$ sampled sequences across different language pairs and parameter settings. Arithmetic sampling also exhibited a lower standard deviation compared to ancestral sampling. The superior performance of arithmetic sampling with multi-sample inference based decoding methods like self-consistency and MBR decoding is due to its ability to generate more diverse samples, enhancing LLM performance in tasks like chain-of-thought reasoning and machine translation without any computational overhead.
@@ A.2 Machine Translation with MBR
font=Huge,labelfont=Huge font=Huge,labelfont=Huge